Create an example probability distribution table to the right in which the expected value is 5. (You must have at least 3 disfferent values for x).
Answer #1

1) X follows binomial ( 10, .5)

    E(X) =5

2) X follows poisson( lamda= 5 )

    E(X) = 5

3) X follows geometric ( q^x * p) , x=0,1,2....

   where q= 5/6, p= 1/6

    E(X) = 5